# CERN-TH-2024-065 ET-0343A-24

### v1

@@ III.2 Gravitational wave generation from a first order phase transition
where $g_{\mu\nu}$ is the background FLRW metric $\mathrm{d}s^{2}=-\mathrm{d}t^{2}+a^{2}(t)\mathrm{d}\mathbf{x}^{2}$, $w$, $p$ are the enthalpy density and pressure of the fluid, $u_{\mu}=a\gamma(1,\mathbf{v})$ is the fluid four-velocity. [^7] Note that we account for the presence of bulk motion in the fluid, which is therefore out of equilibrium. In principle, we should therefore include scalar and vector perturbations in the FLRW metric. However, we neglect their presence, and only focus on the tensor (GW) perturbations generated by the fluid velocity and by the scalar field. This is because the scalar and vector metric perturbations generated by the PT dynamics would be at too small scales to bear any observable significance today. Upon definitions (24)-(26), one obtains for the energy density $\rho_{\mathrm{tot}}=T_{00}$ Athron et al. (2024):
@@
  • PTs with $\alpha\gtrsim\mathcal{O}(1)$ are classified as strong: the vacuum energy is equal or larger than the relativistic fluid energy, and the PT is characterised by supercooling. The vacuum energy leads to a phase of accelerated expansion and, after tunneling, it gets converted mostly into kinetic energy of the bubbles, the thermal component being subdominant. What happens at this point depends on the balance between the vacuum energy and the friction on the bubble wall, due to the interaction with the fluid particles Jinno and Takimoto (2019); Ellis et al. (2020); Lewicki and Vaskonen (2023). In the EWPT, for example, the masses of the fluid particles depend on the value of the Higgs field undergoing the PT. The energy momentum tensors of the scalar field and the surrounding fluid are therefore coupled via an effective friction term, representing the fact that the gradient of the scalar field drives the fluid out of equilibrium by changing the particle masses Ignatius et al. (1994); Kurki-Suonio and Laine (1996a, b); Espinosa et al. (2010). The analysis of the out-of-equilibrium dynamics of this coupled system allows in principle to determine the bubble wall speed $v_{w}$ on a model-by-model basis (as it depends on the specific particle content and interactions of the theory) Moore and Prokopec (1995a, b); Bodeker and Moore (2009); Konstandin et al. (2014); Kozaczuk (2015); Höche et al. (2021a). This is a rather difficult endeavour. Schematically, if the friction is constant (e.g., $\Delta P_{\mathrm{LO}}\sim\Delta m^{2}T^{2}/24$ at linear order for the EW theory, where $\Delta m^{2}$ is the difference in masses of the particles that become heavy during the phase transition), then, whenever the false vacuum energy is larger than the “negative” pressure exerted on the bubble wall by the friction (i.e. $V_{0}(\varphi_{f})>\Delta P_{\mathrm{LO}}$), the bubble wall continuously accelerates until $v_{w}\rightarrow 1$ Moore and Prokopec (1995a, b); Bodeker and Moore (2009). In this so-called runaway regime, very little kinetic energy is transferred to the fluid, and the SGWB is mainly sourced by bubble collision, from the $\Pi_{ij}^{\varphi}=(\partial_{i}\varphi\partial_{j}\varphi)^{TT}$ component of Eq. 51 Ellis et al. (2019b, 2020). However, next-to-leading order processes such as particle splitting at the bubble wall can induce friction terms increasing with $\gamma_{w}=1/\sqrt{1-v_{w}}$ (both behaviours $\Delta P_{\mathrm{NLO}}\propto\gamma_{w}$ Bodeker and Moore (2017); Azatov and Vanvlasselaer (2021); Gouttenoire et al. (2022) and $\Delta P_{\mathrm{NLO}}\propto\gamma_{w}^{2}$ Höche et al. (2021b); Barroso Mancha et al. (2021) have been proposed). In this case, there exists a $v_{w}$ for which the total pressure on the bubble wall balances to zero, $V_{0}(\varphi_{f})-\Delta P_{\mathrm{LO}}-\Delta P_{\mathrm{NLO}}(\gamma_{w})=0$. As soon as the wall reaches this terminal velocity, it stops accelerating. If this occurs before collision, the majority of the released energy is transferred to fluid motion Jinno and Takimoto (2019); Jinno et al. (2019a); Ellis et al. (2019b, 2020); Lewicki and Vaskonen (2023). In a strong PT, the fluid motion takes the form of thin shells of relativistic velocity around the bubbles walls Kamionkowski et al. (1994). The SGWB is therefore mainly sourced by the fluid kinetic energy, from the $\Pi_{ij}^{v}=(w\gamma v_{i}v_{j})^{TT}$ component of Eq. 51, which develops due to the collision of the fluid shells. In general, if the radiation component can be neglected, one has $\kappa_{v}=1-\kappa_{\varphi}$. Furthermore, since $\alpha\gtrsim\mathcal{O}(1)$, Eq. 56 implies $K\simeq 1$.
@@
    Several spectral shapes with distinct features have been proposed in the literature, derived from different approaches to the problem of modelling the anisotropic stresses: e.g., using lattice simulations of the scalar field only Cutting et al. (2018, 2021), of the thin fluid shells surrounding the bubbles Konstandin (2018); Lewicki and Vaskonen (2023), or using analytical models Jinno and Takimoto (2017, 2019). In the rest of this work, we will follow Caprini et al. (2024) and take as placeholder for the SGWB spectrum today, for strong PTs, the result given in Lewicki and Vaskonen (2023), obtained with numerical simulations modeling the contribution from the thin fluid shells. Similarly to Eq. 59, this can be written as a broken power law using the generic template Caprini et al. (2024):
@@
    where the peak of the spectrum in general is $f_{p}=f_{b}\,(-n_{1}/n_{2})^{{1/a_{1}}}$, and the amplitude at the peak is $\Omega_{p}=\Omega_{b}\left\{\left[\left(-n_{2}/n_{1}\right)^{{n_{1}}/{(n_{1}-n_{2})}}+\left(-n_{1}/n_{2}\right)^{-{n_{2}}/{(n_{1}-n_{2})}}\right]/2\right\}^{-(n_{1}-n_{2})/{a_{1}}}$ Caprini et al. (2024). The analysis of Lewicki and Vaskonen (2023) finds that the amplitude at the peak and the peak frequency are (see Eq. 2)
@@
    Recently, from the analysis of the results of numerical simulations with the “Higgsless” approach Jinno et al. (2021, 2023), it has been proposed that a factor should be inserted in the second equality of Eq. 64, since the efficiency in producing kinetic energy in the bulk fluid motion with respect to the single bubble case is not one Hindmarsh et al. (2015); Jinno et al. (2023): one would have then Caprini et al. (2024)
@@
    The shape of the SGWB power spectrum is represented here by the dimensionless factor $[\mathcal{H}_{*}k^{3}F(k)]$. The analytical studies Hindmarsh (2018); Hindmarsh and Hijazi (2019); Roper Pol et al. (2024), in rough agreement with the simulations results Hindmarsh et al. (2014, 2015, 2017); Cutting et al. (2020); Jinno et al. (2023), find that it is proportional to $H_{*}R_{*}/c_{s}$, and that it has the structure of a double broken power law, where both the bubble size $R_{*}$ and the sound shell thickness $\Delta_{w}=|v_{w}-c_{s}|/v_{w}$ are imprinted. We provide below a formula for the SGWB power spectrum from sound waves, taken from Caprini et al. (2024), which can be applied to both weak and intermediate PTs. Note, however, that this formula is preliminary, since it is largely simplified with respect to what found in Gowling and Hindmarsh (2021); Boileau et al. (2023); Gowling et al. (2023), and it is still the subject of ongoing evaluations Roper Pol et al. (2024); Sharma et al. (2023).
@@
    In the meantime, the common approach is to assume that the SGWB spectrum is given by the sum of the two contributions, sound waves and MHD turbulence (inserted in the initial conditions). The kinetic energy in the form of MHD turbulence is assumed to be a fraction of the total bulk kinetic energy. An unknown parameter $\epsilon$ is inserted to represent this fraction Caprini et al. (2016); Roper Pol et al. (2023); Caprini et al. (2024). One further assumes that MHD turbulence is fully developed, so that equipartition has been reached between the kinetic and magnetic turbulent energies. The total energy fraction in the form of MHD turbulence becomes then
@@
    where $\Omega_{\mathrm{kin}}$ and $\Omega_{\mathrm{mag}}$ are the kinetic and magnetic turbulent energy density fractions, and $K$ is given in Eq. 63. The state of the art for SGWB spectra from both sound waves and turbulence is that they can be modelled as double broken power laws. Ref. Caprini et al. (2024) provides a generic template, which can be adapted to both sources, as follows:
@@
    where $K$ is given by Eqs. (64), (66) and (67), and the source duration has been set to the time of formation of non-linarities $\delta\tau_{\mathrm{fin}}\simeq\tau_{\mathrm{nl}}\simeq(3/4)\mathcal{R}_{*}/\sqrt{K}$ (see discussion in the weak PT case), providing the general factor $({H}_{*}R_{*})^{2}/\sqrt{K}/(4/3+{H}_{*}R_{*}/\sqrt{K})$, where we have further incorporated the ${H}_{*}R_{*}$ factor coming from the spectral shape part $[\mathcal{H}_{*}k^{3}F(k)]$. In our analysis we will follow Caprini et al. (2024), which adopts the spectral characteristics found in Jinno et al. (2023) (see, however, Gowling and Hindmarsh (2021); Boileau et al. (2023); Gowling et al. (2023); Roper Pol et al. (2024); Sharma et al. (2023)): we therefore incorporate the factor $0.6$ given in Eq. 68, and set $A_{\text{sw}}\simeq 0.11$ Jinno et al. (2023). Furthermore, the set the frequency breaks $f_{1}$ and $f_{2}$ in Eq. 71 to the two characteristic length scales of the sound wave source, namely the the bubble size $R_{*}$ and the sound shell thickness $\Delta_{w}=|v_{w}-c_{s}|/v_{w}$ Hindmarsh (2018); Hindmarsh and Hijazi (2019); Jinno et al. (2023)
@@
    The slopes are set to $n_{1}=3$, $n_{2}=1$, $n_{3}=-3$, and the form factors $a_{1}=2$, $a_{2}=4$ Jinno et al. (2023). The amplitude at $f_{2}$ is then related to $\Omega_{\mathrm{int}}^{\mathrm{sw}}$ via Caprini et al. (2024)
@@
    The SGWB produced by fully developed, instantaneous MHD turbulence (i.e. inserted in the initial conditions Roper Pol et al. (2022a); Auclair et al. (2022)), instead, is well modeled by the constant-in-time case, Eq. (47). This spectral shape has been validated by (M)HD simulations Roper Pol et al. (2022a); Auclair et al. (2022). We refer to Caprini et al. (2024) for the derivation allowing to put the SGWB spectrum from MHD turbulence derived in Roper Pol et al. (2022a) in the form of Eq. 71, and here we only provide the results. The SGWB spectral amplitude, in terms of the amplitude at the second break $\Omega_{2}$, is
@@
    where $A_{\mathrm{MHD}}\simeq 4.37\times 10^{-3}$. One notes the quadratic behaviour in both the source energy $\Omega_{\mathrm{MHD}}^{2}$ and characteristic scale $(H_{*}R_{*})^{2}$ typical of the constant-in-time scenario, see Eq. 48. Furthermore, the two frequency breaks are also determined within the constant-in-time model. The first break corresponds to the duration of the MHD source, which is of the order of a few eddy turnover times, ${\mathcal{N}}\tau_{\mathrm{e}}\simeq{\mathcal{N}}\mathcal{R}_{*}/(2\pi\sqrt{\varepsilon}\bar{U}_{f})$, with $\mathcal{N}\simeq 2$ Roper Pol et al. (2022a, 2023); Caprini et al. (2024); the second break, instead, corresponds to the characteristic length scale of the MHD turbulence:
@@
    The slopes are $n_{1}=3$, $n_{2}=1$, $n_{3}=-8/3$, and the form factors $a_{1}=4$, $a_{2}\simeq 2.15$ Roper Pol et al. (2022a, 2023); Caprini et al. (2024).
@@ IV.1.4 Constraints from 3G detectors
Figure 2 shows $\Omega_{\rm{GW}}(f)$ for both models, calculated using (87). For Model A (solid, thin line), $G\mu$ has been fixed to $G\mu=7.9\times 10^{-11}$, the value providing the best fit of the EPTA DR2 data set, assuming strings are the unique source of the SGWB detected by PTA, see Antoniadis et al. (2023b). In this case, strings formed with $G\mu=7.9\times 10^{-11}$ would produce a signal in ET/CE with an SNR of more than $500$, providing a wonderful opportunity for coincident detection and thereby a confirmation of the origin of the SGWB in the PTA band. For model B, one the other hand, the value providing the best fit of the EPTA DR2 data set would correspond to $G\mu=2.5\times 10^{-11}$, giving a very loud signal in the Earth-based interferometers band: indeed, such large tensions are already excluded by the LVK data, for which the current constraint on Model B is $G\mu\lesssim{\rm{few}}10^{-15}$ Abbott et al. (2021b). Therefore, for Model B, we have chosen one example of signal detectable only within the Earth-based interferometers band, with $G\mu=4\times 10^{-16}$ (dot-dashed line). [^10] Note that there the spectrum is not flat since we are in the regime $f_{d}\ll f\ll f_{c}$.

### v2

@@ III.2 Gravitational wave generation from a first order phase transition
where $g_{\mu\nu}$ is the background FLRW metric $\mathrm{d}s^{2}=-\mathrm{d}t^{2}+a^{2}(t)\mathrm{d}\mathbf{x}^{2}$, $w$, $p$ are the enthalpy density and pressure of the fluid, $u_{\mu}=\gamma(1,a\mathbf{v})$ is the fluid four-velocity. [^8] Note that we account for the presence of bulk motion in the fluid, which is therefore out of equilibrium. In principle, we should therefore include scalar and vector perturbations in the FLRW metric. However, we neglect their presence, and only focus on the tensor (GW) perturbations generated by the fluid velocity and by the scalar field. This is because the scalar and vector metric perturbations generated by the PT dynamics would be at too small scales to bear any observable significance today. Upon definitions (24)-(26), one obtains for the energy density $\rho_{\mathrm{tot}}=T_{00}$ Athron et al. (2024):
@@
  • PTs with $\alpha\gtrsim\mathcal{O}(1)$ are classified as strong: the vacuum energy is equal or larger than the relativistic fluid energy, and the PT is characterised by supercooling. The vacuum energy leads to a phase of accelerated expansion and, after tunneling, it gets converted mostly into kinetic energy of the bubbles, the thermal component being subdominant. What happens at this point depends on the balance between the vacuum energy and the friction on the bubble wall, due to the interaction with the fluid particles Jinno and Takimoto (2019); Ellis et al. (2020); Lewicki and Vaskonen (2023). In the EWPT, for example, the masses of the fluid particles depend on the value of the Higgs field undergoing the PT. The energy momentum tensors of the scalar field and the surrounding fluid are therefore coupled via an effective friction term, representing the fact that the gradient of the scalar field drives the fluid out of equilibrium by changing the particle masses Ignatius et al. (1994); Kurki-Suonio and Laine (1996a, b); Espinosa et al. (2010). The analysis of the out-of-equilibrium dynamics of this coupled system allows in principle to determine the bubble wall speed $v_{w}$ on a model-by-model basis (as it depends on the specific particle content and interactions of the theory) Moore and Prokopec (1995a, b); Bodeker and Moore (2009); Konstandin et al. (2014); Kozaczuk (2015); Höche et al. (2021a); Barroso Mancha et al. (2021). This is a rather difficult endeavour. Schematically, if the friction is constant (e.g., $\Delta P_{\mathrm{LO}}\sim\Delta m^{2}T^{2}/24$ at linear order for the EW theory, where $\Delta m^{2}$ is the difference in masses of the particles that become heavy during the phase transition), then, whenever the false vacuum energy is larger than the “negative” pressure exerted on the bubble wall by the friction (i.e. $V_{0}(\varphi_{f})>\Delta P_{\mathrm{LO}}$), the bubble wall continuously accelerates until $v_{w}\rightarrow 1$ Moore and Prokopec (1995a, b); Bodeker and Moore (2009). In this so-called runaway regime, very little kinetic energy is transferred to the fluid, and the SGWB is mainly sourced by bubble collision, from the $\Pi_{ij}^{\varphi}=(\partial_{i}\varphi\partial_{j}\varphi)^{TT}$ component of Eq. 51 Ellis et al. (2019b, 2020). However, next-to-leading order processes such as particle splitting at the bubble wall can induce friction terms increasing with $\gamma_{w}=1/\sqrt{1-v_{w}^{2}}$ (both behaviours $\Delta P_{\mathrm{NLO}}\propto\gamma_{w}$ Bodeker and Moore (2017); Azatov and Vanvlasselaer (2021); Gouttenoire et al. (2022); Azatov et al. (2024) and $\Delta P_{\mathrm{NLO}}\propto\gamma_{w}^{2}$ Höche et al. (2021b) have been proposed). In this case, there exists a $v_{w}$ for which the total pressure on the bubble wall balances to zero, $V_{0}(\varphi_{f})-\Delta P_{\mathrm{LO}}-\Delta P_{\mathrm{NLO}}(\gamma_{w})=0$. As soon as the wall reaches this terminal velocity, it stops accelerating. If this occurs before collision, the majority of the released energy is transferred to fluid motion Jinno and Takimoto (2019); Jinno et al. (2019a); Ellis et al. (2019b, 2020); Lewicki and Vaskonen (2023). In a strong PT, the fluid motion takes the form of thin shells of relativistic velocity around the bubbles walls Kamionkowski et al. (1994). The SGWB is therefore mainly sourced by the fluid kinetic energy, from the $\Pi_{ij}^{v}=(w\gamma v_{i}v_{j})^{TT}$ component of Eq. 51, which develops due to the collision of the fluid shells. In general, if the radiation component can be neglected, one has $\kappa_{v}=1-\kappa_{\varphi}$. Furthermore, since $\alpha\gtrsim\mathcal{O}(1)$, Eq. 56 implies $K\simeq 1$.
@@
    Several spectral shapes with distinct features have been proposed in the literature, derived from different approaches to the problem of modelling the anisotropic stresses: e.g., using lattice simulations of the scalar field only Cutting et al. (2018, 2021), of the thin fluid shells surrounding the bubbles Konstandin (2018); Lewicki and Vaskonen (2023), or using analytical models Jinno and Takimoto (2017, 2019). In the rest of this work, we will follow Caprini et al. (2024a) and take as placeholder for the SGWB spectrum today, for strong PTs, the result given in Lewicki and Vaskonen (2023), obtained with numerical simulations modeling the contribution from the thin fluid shells. Similarly to Eq. 59, this can be written as a broken power law using the generic template Caprini et al. (2024a):
@@
    where the peak of the spectrum in general is $f_{p}=f_{b}\,(-n_{1}/n_{2})^{{1/a_{1}}}$, and the amplitude at the peak is $\Omega_{p}=\Omega_{b}\left\{\left[\left(-n_{2}/n_{1}\right)^{{n_{1}}/{(n_{1}-n_{2})}}+\left(-n_{1}/n_{2}\right)^{-{n_{2}}/{(n_{1}-n_{2})}}\right]/2\right\}^{-(n_{1}-n_{2})/{a_{1}}}$ Caprini et al. (2024a). The analysis of Lewicki and Vaskonen (2023) finds that the amplitude at the peak and the peak frequency are (see Eq. 2)
@@
    Recently, from the analysis of the results of numerical simulations with the “Higgsless” approach Jinno et al. (2021, 2023), it has been proposed that a factor should be inserted in the second equality of Eq. 64, since the efficiency in producing kinetic energy in the bulk fluid motion with respect to the single bubble case is not one Hindmarsh et al. (2015); Jinno et al. (2023): one would have then Caprini et al. (2024a)
@@
    The shape of the SGWB power spectrum is represented here by the dimensionless factor $[\mathcal{H}_{*}k^{3}F(k)]$. The analytical studies Hindmarsh (2018); Hindmarsh and Hijazi (2019); Roper Pol et al. (2024), in rough agreement with the simulations results Hindmarsh et al. (2014, 2015, 2017); Cutting et al. (2020); Jinno et al. (2023), find that it is proportional to $H_{*}R_{*}/c_{s}$, and that it has the structure of a double broken power law, where both the bubble size $R_{*}$ and the sound shell thickness $\Delta_{w}=|v_{w}-c_{s}|/v_{w}$ are imprinted. We provide below a formula for the SGWB power spectrum from sound waves, taken from Caprini et al. (2024a), which can be applied to both weak and intermediate PTs. Note, however, that this formula is preliminary, since it is largely simplified with respect to what found in Gowling and Hindmarsh (2021); Boileau et al. (2023); Gowling et al. (2023), and it is still the subject of ongoing evaluations Roper Pol et al. (2024); Sharma et al. (2023).
@@
    In the meantime, the common approach is to assume that the SGWB spectrum is given by the sum of the two contributions, sound waves and MHD turbulence (inserted in the initial conditions). The kinetic energy in the form of MHD turbulence is assumed to be a fraction of the total bulk kinetic energy. An unknown parameter $\epsilon$ is inserted to represent this fraction Caprini et al. (2016); Roper Pol et al. (2023); Caprini et al. (2024a). One further assumes that MHD turbulence is fully developed, so that equipartition has been reached between the kinetic and magnetic turbulent energies. The total energy fraction in the form of MHD turbulence becomes then
@@
    where $\Omega_{\mathrm{kin}}$ and $\Omega_{\mathrm{mag}}$ are the kinetic and magnetic turbulent energy density fractions, and $K$ is given in Eq. 63. The state of the art for SGWB spectra from both sound waves and turbulence is that they can be modelled as double broken power laws. Ref. Caprini et al. (2024a) provides a generic template, which can be adapted to both sources, as follows:
@@
    where $K$ is given by Eqs. (64), (66) and (67), and the source duration has been set to the time of formation of non-linarities $\delta\tau_{\mathrm{fin}}\simeq\tau_{\mathrm{nl}}\simeq(3/4)\mathcal{R}_{*}/\sqrt{K}$ (see discussion in the weak PT case), providing the general factor $({H}_{*}R_{*})^{2}/\sqrt{K}/(4/3+{H}_{*}R_{*}/\sqrt{K})$, where we have further incorporated the ${H}_{*}R_{*}$ factor coming from the spectral shape part $[\mathcal{H}_{*}k^{3}F(k)]$. In our analysis we will follow Caprini et al. (2024a), which adopts the spectral characteristics found in Jinno et al. (2023) (see, however, Gowling and Hindmarsh (2021); Boileau et al. (2023); Gowling et al. (2023); Roper Pol et al. (2024); Sharma et al. (2023)): we therefore incorporate the factor $0.6$ given in Eq. 68, and set $A_{\text{sw}}\simeq 0.11$ Jinno et al. (2023). Furthermore, the set the frequency breaks $f_{1}$ and $f_{2}$ in Eq. 71 to the two characteristic length scales of the sound wave source, namely the the bubble size $R_{*}$ and the sound shell thickness $\Delta_{w}=|v_{w}-c_{s}|/v_{w}$ Hindmarsh (2018); Hindmarsh and Hijazi (2019); Jinno et al. (2023)
@@
    The slopes are set to $n_{1}=3$, $n_{2}=1$, $n_{3}=-3$, and the form factors $a_{1}=2$, $a_{2}=4$ Jinno et al. (2023). The amplitude at $f_{2}$ is then related to $\Omega_{\mathrm{int}}^{\mathrm{sw}}$ via Caprini et al. (2024a)
@@
    The SGWB produced by fully developed, instantaneous MHD turbulence (i.e. inserted in the initial conditions Roper Pol et al. (2022a); Auclair et al. (2022)), instead, is well modeled by the constant-in-time case, Eq. (47). This spectral shape has been validated by (M)HD simulations Roper Pol et al. (2022a); Auclair et al. (2022). We refer to Caprini et al. (2024a) for the derivation allowing to put the SGWB spectrum from MHD turbulence derived in Roper Pol et al. (2022a) in the form of Eq. 71, and here we only provide the results. The SGWB spectral amplitude, in terms of the amplitude at the second break $\Omega_{2}$, is
@@
    where $A_{\mathrm{MHD}}\simeq 4.37\times 10^{-3}$. One notes the quadratic behaviour in both the source energy $\Omega_{\mathrm{MHD}}^{2}$ and characteristic scale $(H_{*}R_{*})^{2}$ typical of the constant-in-time scenario, see Eq. 48. Furthermore, the two frequency breaks are also determined within the constant-in-time model. The first break corresponds to the duration of the MHD source, which is of the order of a few eddy turnover times, ${\mathcal{N}}\tau_{\mathrm{e}}\simeq{\mathcal{N}}\mathcal{R}_{*}/(2\pi\sqrt{\varepsilon}\bar{U}_{f})$, with $\mathcal{N}\simeq 2$ Roper Pol et al. (2022a, 2023); Caprini et al. (2024a); the second break, instead, corresponds to the characteristic length scale of the MHD turbulence:
@@
    The slopes are $n_{1}=3$, $n_{2}=1$, $n_{3}=-8/3$, and the form factors $a_{1}=4$, $a_{2}\simeq 2.15$ Roper Pol et al. (2022a, 2023); Caprini et al. (2024a).
@@ IV.1.4 Constraints from 3G detectors
Figure 2 shows $\Omega_{\rm{GW}}(f)$ for both models, calculated using (87). For Model A (solid, thin line), $G\mu$ has been fixed to $G\mu=7.9\times 10^{-11}$, the value providing the best fit of the EPTA DR2 data set, assuming strings are the unique source of the SGWB detected by PTA, see Antoniadis et al. (2023b). In this case, strings formed with $G\mu=7.9\times 10^{-11}$ would produce a signal in ET/CE with an SNR of more than $500$, providing a wonderful opportunity for coincident detection and thereby a confirmation of the origin of the SGWB in the PTA band. For model B, on the other hand, the value providing the best fit of the EPTA DR2 data set would correspond to $G\mu=2.5\times 10^{-11}$, giving a very loud signal in the Earth-based interferometers band: indeed, such large tensions are already excluded by the LVK data, for which the current constraint on Model B is $G\mu\lesssim{\rm{few}}10^{-15}$ Abbott et al. (2021b). Therefore, for Model B, we have chosen one example of signal detectable only within the Earth-based interferometers band, with $G\mu=4\times 10^{-16}$ (dot-dashed line). [^11] Note that there the spectrum is not flat since we are in the regime $f_{d}\ll f\ll f_{c}$.